OpenStudy (anonymous):

how to find the lateral side surface of the area of a cone... segment y=1/4x from x=0 to x=9 about the x-axis.

OpenStudy (anonymous):

liket his ? 

OpenStudy (anonymous):



OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

wanna use geometry or calcules ?

OpenStudy (anonymous):

Calc

OpenStudy (anonymous):

use the formula for the area of a surface of revolution obtained by rotating the graph curve about the x-axis ,, do u have it ?

OpenStudy (anonymous):

in ur tt book ?

OpenStudy (anonymous):

i guess around y its \(\large 2\pi \int_a^b x.\sqrt {1+f'(x)^2 } dx \) we should make it around y T_T

OpenStudy (anonymous):

wait i ment we need it around x..

OpenStudy (anonymous):

Mhmm!

OpenStudy (anonymous):

So far I have most of it filled in, I'm just missing the dy/dx

OpenStudy (anonymous):

The y=1/4x ?

OpenStudy (anonymous):

And, the dy/dx = 1/2rootx ?

ganeshie8 (ganeshie8):

\[\large SA = \int \limits_0^9 2\pi y ~ ds = \int \limits_0^9 2\pi \left(\frac{1}{4}x\right) \sqrt{1 + \left(\dfrac{dy}{dx}\right)^2} ~dx\]

ganeshie8 (ganeshie8):

\(\large y = \dfrac{1}{4}x \implies \dfrac{dy}{dx} = ?\)